A Bibliography of Publications by, and about, Christian H. Reinsch

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

07 August 2025 Version 1.17

Abstract

This bibliography records publications of Christian H. Reinsch (?? ???? 1934–8 October 2022).

Title word cross-reference

QL [BMRW71]. QR [BMRW71, MRW70, MRW71a, RB68, Rei71c].

12-faces [Rei19a].

2 [Rei14]. 20-Flächer [Rei19b].

464 [Rei73]. 472 [HR73].

507 [HR76].

60 [HR74]. 600 [HR83].

ACM [HR73, Rei73]. action [Rei61]. aids [Rei71a]. Algebra [WR71, Par72]. ALGOL [HR71, HR74]. Algol-60 [HR74]. Algorithm [HR73, Rei73, Rei13, De 89, KA87, MRW70, MRW71a, Rei71c, VBV16, dGR95, Han82, HR76, HR83]. algorithms [BMRW68, BMRW71, DR89]. analysis [DR89, Rei79a, Rei79a]. application [Oli80, SB03]. approximation [Rei74]. Arithmetic [HRR80, Rei79b, RR23]. asynchronen [KS91a]. asynchronous [KS91a]. Automatic [Par72, Rei71a]. automatische [Rei71a]. average [Rei61].

B [Rei14, Rei14]. B-Splines [Rei14, Rei14]. Baker [VBV16]. Balancing [PR69]. band [MRW70, MRW71a]. based [KA87]. Behandlung [Rei79a]. Bernstein [Rei09]. Bernstein-Polynome [Rei09]. best [Rei74]. Book [Par72]. Bulirsch [Mol22a].

C [Par72]. Calculation [HR74, HR71, PR69]. Campbell [VBV16]. case [IP90]. Christian [Ano22, GZ22, Mol22a, Mol22b, Rei22]. Cleve [Mol22a]. Commentaries [CGO07]. Comparison [CES92]. Computation [CGO07, Par72, CH90, Rei71c]. Computer [HRR80, Rei79b, De 89, Rei09, Rei09]. Computer-Typographie [Rei09]. Computing [Gan24, GJ24, Luk80, De 89]. Corner [Mol22a]. cross [Rei61]. cross-section [Rei61]. cubic [DR89]. curves [Rei71a, Rei14]. cyclic [KS91b].

Decomposition [Luk80, ABK18, De 89, GR70, GR07, KS91b, RR23]. differential [Rei61]. differentiellen [Rei61]. distributed [dGR95]. Dodecahedron [Rei19a]. Dodekaeder [Rei19a]. double [RR23]. drawing [Rei71a]. Dreiecke [Rei19b].

E1 [Han82, HR73, HR76]. Effects [Rei77]. Eigenvalues [Rei73, PR69, Rei71c]. eigenvectors [PR69]. einem [KS91a]. Eis [Rei61]. Energieverlustes [Rei61]. energy [Rei61]. equilibrium [Rei19b]. errors [Rei79a]. estimator [SB03]. evaluation [Oli80]. extended [RR23]. extensions [Rei74].

F2 [Rei73]. faces [Rei19a, Rei19b]. Flächer [Rei19a, Rei19b]. Football [Rei16]. Free [Rei14]. Freiform [Rei14]. Freiform-Kurven [Rei14]. Functions [HR74, HR71, Rei67, Rei71a, Rei71b]. Fünfecken [Rei16]. Funktionen [Rei71a]. Fußball [Rei16].

Gene [CGO07]. German [KS91a, Rei61, Rei71a, Rei79a, Rei09, Rei14, Rei16, Rei18, Rei19a, Rei19b]. gleichseitige [Rei19b]. Golub [KS91a, ABK18, CGO07, KS91a]. Greek [Rei19a, Rei19b]. griechisch [Rei19a, Rei19b].

H [CGO07, Par72]. Handbook [Par72]. Hausdorff [VBV16]. Hermitian [Rei71c]. Hilfsmittel [Rei71a]. Householder [MRW68, MRW71b].

ice [Rei61]. icosahedreon [Rei19b]. Ideals [HRR80]. II [Par72, Rei71b]. Ikosaeder [Rei19b]. ILLIAC [Luk80]. Interpolating [HR74, HR71]. Interpolation [Han82, HR73, HR76, HR83, DR89, Rei90]. IV [Luk80].

J [Par72]. Jacobi [De 89].

Kahan [ABK18]. Kurven [Rei14, Rei71a]. kurzes [Rei18].

langsamer [Rei61]. least [GR70, GR07]. like [ABK18]. Linear [Par72, WR71]. logarithmic [Rei61]. logarithmischen [Rei61]. Lösen [Rei18]. loss [Rei61].

machines [dGR95]. Mathematical [Rei71a]. Mathematische [Rei71a]. matrices [BMRW68, BMRW71, DR75, MRW70, MRW71a, RB68]. Matrix [CGO07, Rei73, ABK18, MRW68, MRW71b, PR69, Rei71c]. Measurement [Rei61]. memory [dGR95]. Messung [Rei61]. method [IP90]. Milestones [CGO07]. mittleren [Rei61]. multidimensional [IP90].

Natural [Han82, HR73, HR74, HR76, HR83, HR71]. Neutronen [Rei61]. neutrons [Rei61]. Newbery [Oli80]. Newton [RB68]. numerical [CH90, Rei79a]. numerischen [Rei79a]. NURBs [Rei14].

Oberfläche [Rei16, Rei19b]. Obituary [Mol22b]. one [De 89]. one-sided [De 89]. Oscillation [DR75].

Parallel [KS91b, dGR95]. Parallelisierung [KS91a]. Parallelization [KS91a]. Part [Rei14]. parts [Rei16]. pentagons [Rei16]. Polynome [Rei09]. polynomial [Oli80, Rei09]. Portability [Rei77]. precision [RR23]. Preferences [HRR80, Rei79b]. preserving [DR89, Rei90]. Primer [Rei13]. Principles [HRR80, Rei79b]. problem [IP90]. Procedures [Han82, HR73, HR74, HR76, HR71, HR83]. process [KS91a, KA87]. processor [KS91a]. program [Rei18]. Programm [Rei18]. Prozessorenring [KS91a]. puzzles [Rei18].

QL [BMRW68]. QR [BMRW68]. Quintic [Han82, HR74, HR76, HR83].

Rational [RB68, Rei71c]. Rätseln [Rei18]. Real [Rei73]. Reinsch [IP90, KS91a, Par72, ABK18, Ano22, CES92, GZ22, KS91a, Mol22a, Mol22b, Oli80, Rei22, SB03, VBV16]. Reliable [CH90]. Remark [Han82]. Remez [Rei13]. Review [Par72]. ring [KS91a]. Roland [Mol22a]. rounding [Rei79a]. Rundungsfehlern [Rei79a].

Sard [Rei74]. scattering [Rei61]. scheme [Oli80]. Schoenberg [Rei74]. Sechs [Rei16]. Sechs- [Rei16]. section [Rei61]. seiner [Rei16]. Selected [CGO07]. series [VBV16]. shape [DR89, Rei90, Rei14]. shape-preserving [DR89]. shift [RB68]. short [Rei18]. Side [Rei77]. sided [De 89]. Simplifying [VBV16]. Singular [GR70, GR07, Luk80, RR23, De 89, KS91b]. Singular-Value [Luk80]. six [Rei16]. six- [Rei16]. slow [Rei61]. Smoothing [Rei67, Rei71b, DR75, IP90, KA87]. Software [Rei90]. solution [Rei18]. solutions [GR70, GR07]. solving [IP90]. Some [Rei77]. Speckman [CES92]. Spline [Han82, HR73, HR74, HR76, HR83, DR89, DR75, HR71, KA87, Rei67, Rei71b, Rei90, SB03]. Splines [CES92, Rei14]. squares [GR70, GR07]. stable [Rei71c]. stochastic [KA87]. storing [KS91b]. Streuung [Rei61]. Striving [Rei77]. SUDOKU [Rei18]. SUDOKU-Rätseln [Rei18]. surface [Rei19b]. surfaces [Rei16]. SVD [ABK18, Gan24, GJ24, Mol22a]. SVD-like [ABK18]. Symmetric [Rei73, BMRW68, BMRW71, MRW68, MRW70, MRW71a, MRW71b, RB68]. Symplectic [ABK18].

Talbot [dGR95]. Teil [Rei14]. Teilen [Rei16]. theory [Rei74]. transformation [Oli80, RB68]. Translation [HR83]. Treatment [Rei79a]. triangles [Rei19b]. Tridiagonal [Rei73, RB68, Rei71c]. tridiagonalization [MRW68, MRW71b]. Two [Rei74, DR89]. Typographie [Rei09]. typography [Rei09].

Value [Luk80, De 89, GR70, GR07, KS91b, RR23]. vector [De 89]. Verfahrens [KS91a]. Vol [Par72].

Wasser [Rei61]. water [Rei61]. Wilkinson [Par72]. Wirkungsquerschnittes [Rei61]. Works [CGO07].

Zeichnen [Rei71a]. zum [Rei18]. Zwölf [Rei19a]. Zwölf-Flächer [Rei19a].

References

Agoujil:2018:SGK

[ABK18] S. Agoujil, A. H. Bentbib, and A. Kanber. Symplectic Golub– Kahan–Reinsch SVD-like matrix decomposition. Journal of Advanced Mathematical Studies, 11(2):203–213, 2018. ISSN 2065-3506,2065-5851.

Anonymous:2022:CR

[Ano22] Anonymous. Christian Reinsch. Wikipedia article, 2022. URL

          https://en.wikipedia.org/wiki/Christian_Reinsch.

Bowdler:1968:QQA

[BMRW68] H. Bowdler, R. S. Martin, C. Reinsch, and J. H. Wilkinson. The QR and QL algorithms for symmetric matrices. Numerische Mathematik, 11:293–306, 1968. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic). URL https://eudml.org/doc/

131827.

Bowdler:1971:ASM

[BMRW71] H. Bowdler, R. S. Martin, C. H. Reinsch, and J. H. Wilkinson. The QR and QL algorithms for symmetric matrices. In Wilkinson and Reinsch [WR71], pages 227–240. ISBN 0-387-05414-6, 3-540-05414-6. ISSN 0072-7830. LCCN QA251 .W67.

Carter:1992:CRS

[CES92] C. K. Carter, G. K. Eagleson, and B. W. Silverman. A comparison of the Reinsch and Speckman splines. Biometrika, 79(1):81–91, March 1992. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2337149.

Chan:2007:MMC

[CGO07] Raymond H. Chan, Chen Greif, and Dianne P. O’Leary, editors. Milestones in Matrix Computation: the Selected Works of Gene H. Golub with Commentaries. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920681-3. xi + 565 + 3

pp. LCCN QA188 .G67 2007. URL http://www.loc.gov/catdir/
enhancements/fy0737/2007276086-d.html.

Cox:1990:RNC

[CH90] M. G. Cox and S. Hammarling, editors. Reliable numerical computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

DeRijk:1989:OSJ

[De 89] P. P. M. De Rijk. A one-sided Jacobi algorithm for computing the singular value decomposition on a vector computer. SIAM Journal on Scientific and Statistical Computing, 10(2):359–371, March 1989. CODEN SIJCD4. ISSN 0196-5204.

deRosa:1995:PTA

[dGR95] M. A. de Rosa, G. Giunta, and M. Rizzardi. Parallel Talbot’s algorithm for distributed memory machines. Parallel Computing, 21 (5):783–801 (or 783–802??), May 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.com/cgi-bin/cas/tree/store/parco/cas_sub/
browse/browse.cgi?year=1995&volume=21&issue=5&aid=974.

Demmler:1975:OMS

[DR75] A. Demmler and C. Reinsch. Oscillation matrices with spline smoothing. Numerische Mathematik, 24(5):375–382, 1975. CO-DEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

URL https://eudml.org/doc/132350.

Dauner:1989:ATA

[DR89] Herbert Dauner and Christian H. Reinsch. An analysis of two algorithms for shape-preserving cubic spline interpolation. IMA Journal of Numerical Analysis, 9(3):299–314, 1989. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Gander:2024:CSa

[Gan24] Walter Gander. Computing the SVD. Netlib Web site., July

5, 2024. URL https://na-digest.coecis.cornell.edu/na-
digest-html/24/v24n10.html#1.

Gander:2024:CSb

[GJ24] Walter Gander and Johann Joss. Computing the SVD. Report, [retired from ETH], Zurich, Switzerland, November 24, 2024. 25

pp. URL https://people.inf.ethz.ch/gander/talks/SVDnew.
pdf.

Golub:1970:SVD

[GR70] G. H. Golub and C. Reinsch. Singular value decomposition and least squares solutions. Numerische Mathematik, 14:403–420, 1970. CO-DEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

URL https://eudml.org/doc/131963.

Golub:2007:SVD

[GR07] Gene H. Golub and Christian H. Reinsch. Singular value decomposition and least squares solutions. In Chan et al. [CGO07], pages 160–180. ISBN 0-19-920681-3. LCCN QA188 .G67

2007. URL http://www.loc.gov/catdir/enhancements/fy0737/
2007276086-d.html.

Gander:2022:CR

[GZ22] Walter Gander and Christoph Zenger. Christian Reinsch. Netlib Web site, October 12, 2022. URL https://listserv.

utk.edu/cgi-bin/wa?A2=NADIGEST;705e79c0.2210c;   https:
//netlib.org/na-digest-html/22/v22n34.html#1.

Hanson:1982:RPQ

[Han82] R. J. Hanson. Remark on “Algorithm 507: Procedures for quintic natural spline interpolation [E1]”. ACM Transactions on Mathematical Software, 8(3):334, September 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HR76].

Herriot:1971:APC

[HR71] John G. Herriot and Christian H. Reinsch. ALGOL 60 procedures for the calculation of interpolating natural spline functions, volume 71-200 of Rep., Dep. Comput. Sci., Stanford Univ. Stanford, CA: Computer Science Department, School of Humanities and Sciences, Stanford University, 1971.

Herriot:1973:AAP

[HR73] John G. Herriot and Christian H. Reinsch. ACM Algorithm 472: Procedures for natural spline interpolation [E1]. Communications of the ACM, 16(12):763–768, December 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herriot:1974:APC

[HR74] John G. Herriot and Christian H. Reinsch. Algol-60 procedures for the calculation of interpolating natural quintic spline functions. Report, Department of Computer Science, Stanford University, Stanford, CA, January 1974.

Herriot:1976:APQ

[HR76] John G. Herriot and Christian H. Reinsch. Algorithm 507: Procedures for quintic natural spline interpolation [E1]. ACM Transactions on Mathematical Software, 2(3):281–289, September 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Han82].

Herriot:1983:ATA

[HR83] John G. Herriot and Christian H. Reinsch. Algorithm 600: Translation of Algorithm 507: Procedures for quintic natural spline interpolation. ACM Transactions on Mathematical Software, 9(2): 258–259, June 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hull:1980:PPI

[HRR80] Thomas E. Hull, Christian H. Reinsch, and John R. Rice. Principles, preferences and ideals for computer arithmetic. Technical report TR-339, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1, 1980.

13 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1980/TR%2080-339.pdf.

Ignatov:1990:RMS

[IP90] M. I. Ignatov and A. B. Pevnyĭ. Reinsch’s method for solving the smoothing problem in the multidimensional case. Zhurnal vychislitel’noi matematiki i matematicheskoi fiziki, 30(2):186–192, 334, 1990. CODEN ZVMFAN. ISSN 0044-4669.

Kohn:1987:NAS

[KA87] Robert Kohn and Craig F. Ansley. A new algorithm for spline smoothing based on smoothing a stochastic process. SIAM Journal on Scientific and Statistical Computing, 8(1):33–48, January 1987. CODEN SIJCD4. ISSN 0196-5204.

Kockler:1991:PGR

[KS91a] N. Köckler and M. Simon. Die Parallelisierung des Golub–Reinsch– Verfahrens auf einem asynchronen Prozessorenring. (German) [Parallelization of the Golub–Reinsch process on an asynchronous processor ring]. Zeitschrift für Angewandte Mathematik und Mechanik, 71(6):T803–T806, ???? 1991. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic). Bericht über die Wissenschaftliche Jahrestagung der GAMM (Hannover, 1990).

Kockler:1991:PSV

[KS91b] N. Kockler and M. Simon. Parallel singular value decomposition with cyclic storing. Parallel Computing, 17(1):39–47, April 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Luk:1980:CSV

[Luk80] Franklin T. Luk. Computing the singular-value decomposition on the ILLIAC IV. ACM Transactions on Mathematical Software, 6 (4):524–539, December 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Moler:2022:CCC

[Mol22a] Cleve Moler. Cleve’s corner: Christian Reinsch, Roland Bulirsch, and the SVD. MathWorks Web site., October 23, 2022. URL

http://feeds.feedburner.com/mathworks/moler;   https://
blogs.mathworks.com/cleve/2022/10/23/christian-reinsch-
roland-bulirsch-and-the-svd/.

Moler:2022:OCR

[Mol22b] Cleve Moler. Obituary: Christian Reinsch. SIAM News Web site.,

October 28, 2022. URL https://www.siam.org/publications/
siam-news/articles/obituary-christian-reinsch/.

Martin:1968:HTS

[MRW68] R. S. Martin, C. Reinsch, and J. H. Wilkinson. Householder’s tridiagonalization of a symmetric matrix. Numerische Mathematik, 11: 181–195, 1968. CODEN NUMMA7. ISSN 0029-599x (print), 0945-

3245 (electronic). URL https://eudml.org/doc/131819.

Martin:1970:ABS

[MRW70] R. S. Martin, C. Reinsch, and J. H. Wilkinson. The QR algorithm for band symmetric matrices. Numerische Mathematik, 16:85–92, 1970. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (elec-

tronic). URL https://eudml.org/doc/132019.

Martin:1971:ABS

[MRW71a] R. S. Martin, C. H. Reinsch, and J. H. Wilkinson. The QR algorithm for band symmetric matrices. In Wilkinson and Reinsch [WR71], pages 266–272. ISBN 0-387-05414-6, 3-540-05414-6. ISSN 0072-7830. LCCN QA251 .W67.

Martin:1971:HTS

[MRW71b] R. S. Martin, C. H. Reinsch, and J. H. Wilkinson. Householder’s tridiagonalization of a symmetric matrix. In Wilkinson and Reinsch [WR71], chapter 13, pages 212–226. ISBN 0-387-05414-6, 3-540-05414-6. ISSN 0072-7830. LCCN QA251 .W67.

Oliver:1980:ANT

[Oli80] J. Oliver. On the application of Newbery’s transformation to the Reinsch polynomial evaluation scheme. Journal of Computational and Applied Mathematics, 6(1):43–52, March 1980. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0771050X80900169.

Partlett:1972:BRB

[Par72] B. Partlett. Book review: Handbook for Automatic Computation, Vol. II, Linear Algebra (J. H. Wilkinson and C. Reinsch). SIAM Review, 14(4):658–661, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1969:BMC

[PR69] B. N. Parlett and C. Reinsch. Balancing a matrix for calculation of eigenvalues and eigenvectors. Numerische Mathematik, 13:293– 304, 1969. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245

(electronic). URL https://eudml.org/doc/131920.

Reinsch:1968:RTN

[RB68] C. Reinsch and F. L. Bauer. Rational QR transformation with Newton shift for symmetric tridiagonal matrices. Numerische Mathematik, 11:264–272, 1968. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic). URL https://eudml.org/doc/

131824.

Reinsch:1961:MDW

[Rei61] Christian H. Reinsch. Messung des differentiellen Wirkungsquerschnittes und des mittleren logarithmischen Energieverlustes bei der Streuung langsamer Neutronen an Wasser und Eis. (German) [Measurement of the differential cross-section of action and the average logarithmic energy loss when scattering slow neutrons on water and ice]. Doctor rerum naturalium (Dr. rer. nat.) thesis, Technische Universität München, Munich, Germany, 1961. URL https:

//dl.acm.org/profile/81100396422.

Reinsch:1967:SSF

[Rei67] Christian H. Reinsch. Smoothing by spline functions. Numerische Mathematik, 10(3):177–183, October 1967. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

URL https://eudml.org/doc/131782;   https://tlakoba.w3.
uvm.edu/AppliedUGMath/auxpaper_Reinsch_1967.pdf. See Part

II [Rei71b].

Reinsch:1971:MHA

[Rei71a] Christian H. Reinsch. Mathematische Hilfsmittel für das automatische Zeichnen von Funktionen und Kurven. (German) [Mathematical aids for the automatic drawing of functions and curves]. Fachtagung Computer Graphics, Ges. Informatik, Bericht 2, 309–338 (1971)., 1971.

Reinsch:1971:SSF

[Rei71b] Christian H. Reinsch. Smoothing by spline functions. II. Numerische Mathematik, 16(5):451–454, February 1971. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic). URL

https://eudml.org/doc/132051. See Part I [Rei67].

Reinsch:1971:SRA

[Rei71c] Christian H. Reinsch. A stable, rational QR algorithm for the computation of the eigenvalues of an Hermitian, tridiagonal matrix. Mathematics of Computation, 25(115):591–597, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reinsch:1973:AAE

[Rei73] Christian H. Reinsch. ACM Algorithm 464: Eigenvalues of a real, symmetric, tridiagonal matrix [F2]. Communications of the ACM, 16(11):689, November 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reinsch:1974:TES

[Rei74] Christian Reinsch. Two extensions of the Sard–Schoenberg theory of best approximation. SIAM Journal on Numerical Analysis, 11: 45–51, 1974. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reinsch:1977:SSE

[Rei77] Christian Reinsch. Some side effects of striving for portability. In Wayne Cowell, editor, Portability of Numerical Software, Workshop, Oak Brook, Illinois, June 21–23, 1976, volume 57 of Lecture Notes in Computer Science, pages 3–4. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 3-540-08446-0.

Reinsch:1979:BRN

[Rei79a] Christian Reinsch. Die Behandlung von Rundungsfehlern in der numerischen Analysis. (German) [Treatment of rounding errors in numerical analysis]. In Jahrbuch Überblicke Mathematik, 1979, pages 43–62. Bibliographisches Institut, Mannheim, Germany, 1979. ISBN 3-411-01559-4.

Reinsch:1979:PPC

[Rei79b] Christian H. Reinsch. Principles and preferences for computer arithmetic. ACM SIGNUM Newsletter, 14(1):12–27, March 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Reinsch:1990:SSP

[Rei90] Christian H. Reinsch. Software for shape preserving spline interpolation. In Cox and Hammarling [CH90], pages 267–276. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (born 1919, died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Reinsch:2009:BPC

[Rei09] Christian Reinsch. Bernstein-Polynome für die Computer-Typographie. (German) [Bernstein polynomial for computer typography]. Class notes, May 19, 2009. URL https:/

/drive.google.com/file/d/1B-LwGCOsuY3xfzamfScBUQBmHo7o-
z64/view?usp=sharing.

Reinsch:2013:PRA

[Rei13] Christian Reinsch. A primer on the Remez algorithm. Class

notes, October 17, 2013. URL https://drive.google.com/file/
d/1a1RKLOTpGJX9n_UoEjqM30knsV0YgSB-/view.

Reinsch:2014:PFK

[Rei14] Christian Reinsch. Teil 2: Freiform-Kurven mit B-Splines und NURBs. (German) [Part 2: Free shape curves with Bsplines and NURBs]. Class notes, January 21, 2014. URL

https://drive.google.com/file/d/1Yoc-Uuoc0EhdracQnBOeR_
s5EMagp6Gb/view?usp=sharing.

Reinsch:2016:FSF

[Rei16] Christian Reinsch. Fußball mit 20 Sechs- und 12 Fünfecken als Teilen seiner Oberfläche. (German) [Football with 20 sixand 12 pentagons as parts of its surfaces]. Class notes, De-

cember 16, 2016. URL https://drive.google.com/open?id=
1yWBqzhpAA5JOrGaV3Xerf8SiROtCW82d.

Reinsch:2018:KPL

[Rei18] Christian Reinsch. Ein kurzes Programm zum Lösen von SUDOKU-Rätseln. (German) [A short program for the solution of Sudoku puzzles]. Class notes, July 29, 2018. URL https://drive.google.

com/file/d/1Ew5DjomerB1Z7BmIkqhKbXhAvan1sL1i/view?usp=sharing.

Reinsch:2019:DGZ

[Rei19a] Christian Reinsch. Den Dodekaeder (griechisch, Zwölf-Flächer). (German) [The dodecahedron (Greek, 12-faces)]. Class notes,

March 26, 2019. URL https://drive.google.com/file/d/
1SldgysrOhyIjb5jKIXrNc75Kvdq6DL5f/view.

Reinsch:2019:IGF

[Rei19b] Christian Reinsch. Der Ikosaeder (griechisch, “20-Flächer;;) hat 20 gleichseitige Dreiecke als Oberfläche. (German) [The icosahedreon (Greek, “20 faces”) has 20 equilibrium triangles on its surface]. Class notes, March 25, 2019. URL https://drive.google.com/

file/d/1YEovz_i4MInUC56tkJifevtmVYjP0iIi/view.

Reinsch:2022:CR

[Rei22] Christian Reinsch. Christian Reinsch. Web site, 2022. URL https:

//sites.google.com/view/christianreinsch.

Reinsch:2023:SVD

[RR23] Christian Reinsch and Mathias Richter. Singular value decomposition in extended double precision arithmetic. Numerical Algorithms, 93(3):1137–1155, 2023. CODEN NUALEG. ISSN 1017-1398,1572-9265.

Sanusi:2003:RSE

[SB03] Wahidah Sanusi and I. Nyoman Budiantara. The Reinsch spline estimator and its application. Journal of the Indonesian Mathematical Society, 9(1):1–8, 2003. ISSN 0854-1388.

Van-Brunt:2016:SRA

[VBV16] Alexander Van-Brunt and Matt Visser. Simplifying the Reinsch algorithm for the Baker–Campbell–Hausdorff series. Journal of Mathematical Physics, 57(2):023507, 19, February 2016. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation, Editors: F. L. Bauer, A. S. Householder, F. W. J. Olver, H. Rutishauser, K. Samelson and E. Stiefel. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6. ISSN 0072-7830. viii + 439 pp. LCCN QA251 .W67.